OpenStudy (anonymous):

Find the product of: 3 1 10 - x 11- 4 2

OpenStudy (anonymous):

\[10\frac{3}{4}=\frac{10\times4+3}{4}\] \[11\frac{1}{2}=\frac{11\times2+1}{2}\] can you multiply yourself now? please please please

OpenStudy (diyadiya):

\[10\frac{3}{4} = \frac{10 * 4 + 3}{4} = \frac{43}{4}\]\[11 \frac{1}{2}= \frac{11*2+1}{2}= \frac{23}{2}\] \[\frac{43}{4} \times \frac{23}{2}=\frac{989}{8}\]\[123 \frac{5}{8}\]